# Bayesian Empirical Bayes: Simultaneous Inference from Probabilistic Symmetries

### v1

@@ Abstract
Contact: bw2766@columbia.edu, ew2760@columbia.edu, david.blei@columbia.edu
@@ 2.2 Examples of $\Phi$-invariant priors
$u_{i},v_{j},u_{ij}$ $\stackrel{{\scriptstyle iid}}{{\sim}}\textrm{unif}(0,1)\quad z_{ij}=g_{\mathrm{sep}}(u_{i},v_{j},u_{ij})\quad\text{for all }i,j\in\mathbb{N}.$ (17)
@@ 2.3 The finite population model
$\bm{z}_{n}$ $\sim\mathrm{p}(\bm{z}_{n}\,|\,\textrm{g})$ (20)
$x_{\omega}$ $\sim\mathrm{p}(x\,|\,z_{\omega})\quad\omega\in S_{n}.$ (21)
@@ 2.4 Bayesian empirical Bayes
where both these terms are known. The only unknown component in Eq.˜22 is $\mathrm{p}^{\star}(\textrm{g}\,|\,\bm{x}_{n})$, which depends on the unknown $\mu^{\star}$.
@@
This approximation is motivated by both asymptotic and computational considerations. The observed dataset $\bm{x}_{n}$ is generated from a single draw $\textrm{g}^{\star}$. So asymptotically, we expect the ideal posterior $\mathrm{p}^{\star}(\textrm{g}\,|\,\bm{x}_{n})$ to concentrate at $\textrm{g}^{\star}$ as $n\to\infty$. In particular, by Theorem˜1, the law $\mathrm{p}(\bm{z}\,|\,\textrm{g})$ is $\Phi$-ergodic and hence identifiable (see, e.g., (Kallenberg1997foundations, Lemma 27.1) and (Lindenstrauss2001, Theorem 1.2)). Under a characteristic likelihood, such as the normal model with known variance, distinct priors $\mathrm{p}(\bm{z}\,|\,\textrm{g})$ induce distinct marginals $\mathrm{p}(\bm{x}\,|\,\textrm{g})$, so g is identifiable from $\mathrm{p}(\bm{x}\,|\,\textrm{g})$. Posterior concentration then follows from Doob’s theorem (Ghosal2017, Ch. 6). Computationally, although the marginal likelihood in Eq.˜24 may be intractable, we can use variational inference to approximate it (Blei2017).
Finally, we substitute the point-estimate approximation of the prior into the posterior of Eq.˜22. This gives the approximation
@@
Our procedure is called Bayesian empirical Bayes (BEB). We are given data $\bm{x}_{n}$ and assume a symmetry group $\Phi$ for the latent variables $\bm{z}_{n}$. We use the ergodic representation theorem to form the expanded hierarchical model (Eq.˜19), which includes the invariant variable g and the corresponding ergodic family $\mathrm{p}(\bm{z}\,|\,\textrm{g})$. We parameterize g and estimate it from data via maximum marginal likelihood (Eq.˜24) or a suitable surrogate. Finally, we plug the estimate $\hat{\textrm{g}}$ into the conditional posterior (Eq.˜25). Algorithm˜1 presents this recipe.
@@ 3 New Models for Empirical Bayes
We now use BEB to design new methods for empirical Bayes inference. In our first example, we use BEB to recover the classical empirical Bayes g-modeling for the normal means model (§˜3.1). This recovers classical EB methods (Efron2015deconvolution; Soloff2021). We then extend EB to three additional probabilistic symmetries: separate/joint exchangeability in arrays (§˜3.2), stationarity in spatial data (§˜3.3), and relative exchangeability conditional on covariates (§˜D.3). In each case, we follow the recipe of Algorithm˜1. We start with a dataset, a symmetry group, and a likelihood. We then design an empirical Bayes model–tailored to the symmetry–and approximate the posterior of the latent structured variables.
@@ 3.1 Classical Example Revisited: Normal Means Model
We assume the true prior $\mathrm{p}^{\star}(\bm{z})$ is exchangeable. By de Finetti’s theorem, we can write Eq.˜26 as the following generative process:
@@
Here $g_{\rm perm}$ is a random prior over $z$. In short, de Finetti’s representation specializes Theorem˜1 to sequences, with $\prod_{i=1}^{\infty}\textrm{g}_{\rm\textrm{perm}}(z_{i})$ the ergodic distribution.
@@
In Bayesian EB, we approximate $\mathrm{p}^{\star}(\textrm{g}_{\rm perm}\mid\bm{x}_{n})$ with its MMLE,
@@
If we restrict $g_{\rm perm}$ to a parametric family, then this recovers parametric empirical Bayes estimation, or type-II maximum likelihood. If we consider all possible priors, then it recovers the nonparametric maximum likelihood estimation (NPMLE) approach.
With $\hat{g}_{\rm perm}$ in hand, we approximate the target posterior from Eq.˜28,
@@
The goal is to recover the clean image $\bm{z}^{\star}_{n,n}$ from the observed noisy image $\bm{x}_{n,n}$. The maximum likelihood estimator is the simply the noisy image which has a MSE of $10.291$. The exchangeable sequence approach treats the problem as a simple normal means problem with known variance $\sigma_{0}^{2}$ and applies the NPMLE procedure. The resulting MSE is 3.613.61–a substantial reduction from the error of the observed image.
@@ 3.2 Bayesian empirical Bayes for matrix recovery
$u_{i},v_{j},u_{ij}$ $\stackrel{{\scriptstyle iid}}{{\sim}}\textrm{unif}[0,1],\quad\quad\quad i\in[n],j\in[n]$ (32)
$z_{ij}$ $=\textrm{g}_{\textrm{sep}}(u_{i},v_{j},u_{ij}),\quad$ (33)
$x_{ij}$ $\sim\mathcal{N}(z_{ij},\,\tau_{ij}^{-1}).$ (34)
@@
Exact evaluation of Eq.˜35 is intractable because the integral is high–dimensional. We approximate it by maximizing an evidence lower bound (ELBO) under a structured variational family. The estimate $\hat{\textrm{g}}_{\textrm{sep},\theta}$ is the variational MMLE. See Appendix˜D for details.
@@
If the array is jointly exchangeable then it is invariant when rows and columns are permuted together. In this case, Eq.˜30 simplifies to
@@
If the array is relatively exchangeable conditioned on row covariates $\bm{y}_{n}=(y_{i})_{i=1}^{n}$ with $y_{i}\in\mathcal{Y}$ and column covariates $\bm{a}_{p}=(a_{j})_{j=1}^{p}$ with $a_{j}\in\mathcal{A}$, then Eq.˜30 expands to
@@
We revisit the example of denoising an MNIST "5" from Figure˜1. The image is treated as a noisy matrix and we apply the empirical Bayes matrix recovery under separate and joint exchangeability. We report $100\times\textrm{MSE}$ between the posterior means $\hat{\bm{z}}_{n,n}$ and the true pixel value $\bm{z}^{\star}_{n,n}$. The fourth and fifth panels in Figure˜1 show the results. EBMR with a jointly exchangeable prior (EBMR–Joint) achieves the lowest error, followed by EBMR with a separately exchangeable prior (EBMR–Sep), then EBMR with a fully exchangeable prior (NPMLE). This ordering aligns with the symmetry hierarchy full $\Rightarrow$ separate $\Rightarrow$ joint: weaker (less restrictive) symmetry assumptions enlarge the class of admissible priors and increase flexibility. On the other hand, weaker symmetry assumptions also increase the sample and computational complexity for estimating g. This creates a trade-off. In our setting, joint exchangeability offers a sweet spot between expressiveness and efficiency.
@@ 3.3 Bayesian empirical Bayes for spatial regression
where $\beta\in\mathbb{R}^{p}$ are regression coefficients and $\bm{z}=(z_{\omega})_{\omega\in\Omega}$ is a latent process that captures spatial dependence. We place a flat prior $\pi(\beta)\propto 1$, so the estimator of $\beta$ coincides with weighted least squares. The resulting model along with Eq.˜37 coincide with the standard model in Bayesian spatial regression (handcock1993bayesian; Stein1999; DeOliveira1997spatial; Berger2001spatial; Banerjee2008; Banerjee2014spatial; Li2024spatial).
@@
$g\sim\mu^{\star},\qquad\bm{z}\mid g\sim g,$ (38)
@@
We have defined the invariant variable $\bm{g}$ as the spectral density of a stationary kernel for a Gaussian process. With this definition, the BEB model is
@@
Figure 1: EB spatial regression on a noisy MNIST “5.” Top: original (left), noisy (middle), and NPMLE (right). Bottom: EBMR–sep (left), EBMR–joint (middle), and EB spatial regression (right). Errors are $100\times\textrm{MSE}$ of posterior means; EBMR with a jointly exchangeable prior performs best, followed by EBMR with a separately exchangeable prior and EB spatial regression, then NPMLE.
@@ 4.3 Empirical Bayes Spatial Regression
where the formula for $k_{w,\sigma,\mu}$ is in Eq.˜69. Equivalently, it is a stationary kernel with a three-component Gaussian mixture $0.5\mathcal{N}(0,1^{2})+0.3\mathcal{N}(0.15,0.2^{2})+0.2\mathcal{N}(-0.25,0.35^{2})$ as the spectral density. We draw a single $\bm{z}_{n}^{\star}\sim\mathcal{N}(0,\Sigma^{\star})$ as the ground truth.
@@
Figure˜3 shows the results. EB spatial regression uniformly dominates the regression-adjusted NPMLE, which indicates that stationarity is a more informative symmetry than full exchangeability for this example. At each noise level, the MSE of the EB spatial estimator of $\bm{z}_{n}^{\star}$ decreases sharply with the sample size, while the MSE of the regression-adjusted NPMLE remains essentially flat as $n$ increases. The performance of the two methods is comparable only for small samples (e.g., $n=100$); for larger $n$, the spatial EB estimator achieves substantially lower error. In Figure˜10 of Appendix˜G, we show that the MSE of the posterior-mean estimator for $\beta$ also decreases with $n$ under the spatial EB model, which suggests that the posterior of $\beta$ satisfies posterior consistency.
@@ 5.2 Denoising and Interpolating Air Quality Data in NYC
Figure˜5 shows empirical Bayes posterior means with $95\%$ credible bands. The posterior means shows the shrinkage and pooling effects of EB spatial regression: extreme weekly averages (e.g., mid-2023) are pulled toward the mean. Where observations are dense, the posterior uncertainty around the denoised weekly averages is narrow; where data are missing, it is wider. Within Manhattan (Herald Square, Chinatown, Lower East Side), the model borrows heavily across sites—for example, when Chinatown is entirely missing from mid-2021 to early-2022, the model interpolates from Herald Square and the Lower East Side to produce tight bands; similarly, during mid-2023 to early-2024, the model imputes the Lower East Side using the other two Manhattan sites, which recovers a decreasing, cyclical trend. By contrast, Queens North contributes less to interpolating Manhattan gaps (e.g., mid-2022 to early-2023). The reason for this different shrinking behavior is that the EB model finds two spatio–temporal regimes–one for Manhattan and one for northern Queens; see Figure˜9 in §˜G.3.
@@ D.1 EBMR with a Separately Exchangeable Prior
To approximate the intractable marginal likelihood in Eq.˜35, we posit the following structured variational family
@@
Following Eq.˜49, the empirical Bayes estimate minimizes the negative ELBO over the network and grid weights:
@@ Appendix E Algorithms for Empirical Bayes Spatial Regression
where each $\Sigma_{k}=\mathrm{diag}(\sigma_{k,1}^{2},\ldots,\sigma_{k,d}^{2})$ is diagonal. By Bochner’s theorem, Eq.˜68 induces the kernel
@@ G.2 Covariate-Assisted Empirical Bayes with Model Misspecification
G.3 Additional Results for NYC Air Quality

### v2

@@ Abstract
Contact: bw2766@columbia.edu, enawe@dtu.dk, david.blei@columbia.edu
@@ 2.2 Examples of $\Phi$-invariant priors
$u_{i},v_{j},u_{ij}$ $\stackrel{{\scriptstyle iid}}{{\sim}}\textrm{unif}(0,1)\quad z_{ij}=\textrm{g}_{\mathrm{sep}}(u_{i},v_{j},u_{ij})\quad\text{for all }i,j\in\mathbb{N}.$ (17)
@@ 2.3 The finite population model
$\bm{z}_{n}$ $\sim\mathrm{p}(\bm{z}_{n}\,|\,\textrm{g})$
$x_{\omega}$ $\sim\mathrm{p}(x\,|\,z_{\omega})\quad\omega\in S_{n}.$
@@ 2.4 Bayesian empirical Bayes
where both these terms are known. The only unknown component in Eq.˜20 is $\mathrm{p}^{\star}(\textrm{g}\,|\,\bm{x}_{n})$, which depends on the unknown $\mu^{\star}$.
@@
This approximation is motivated by both asymptotic and computational considerations. The observed dataset $\bm{x}_{n}$ is generated from a single draw $\textrm{g}^{\star}$. So asymptotically, we expect the ideal posterior $\mathrm{p}^{\star}(\textrm{g}\,|\,\bm{x}_{n})$ to concentrate at $\textrm{g}^{\star}$ as $n\to\infty$. In particular, by Theorem˜1, the law $\mathrm{p}(\bm{z}\,|\,\textrm{g})$ is $\Phi$-ergodic and hence identifiable (see, e.g., (Kallenberg1997foundations, Lemma 27.1) and (Lindenstrauss2001, Theorem 1.2)). Under a characteristic likelihood, such as the normal model with known variance, distinct priors $\mathrm{p}(\bm{z}\,|\,\textrm{g})$ induce distinct marginals $\mathrm{p}(\bm{x}\,|\,\textrm{g})$, so g is identifiable from $\mathrm{p}(\bm{x}\,|\,\textrm{g})$. Posterior concentration then follows from Doob’s theorem (Ghosal2017, Ch. 6). Computationally, although the marginal likelihood in Eq.˜22 may be intractable, we can use variational inference to approximate it (Blei2017).
Finally, we substitute the point-estimate approximation of the prior into the posterior of Eq.˜20. This gives the approximation
@@
Our procedure is called Bayesian empirical Bayes (BEB). We are given data $\bm{x}_{n}$ and assume a symmetry group $\Phi$ for the latent variables $\bm{z}_{n}$. We use the ergodic representation theorem to form the expanded hierarchical model (Eq.˜19), which includes the invariant variable g and the corresponding ergodic family $\mathrm{p}(\bm{z}\,|\,\textrm{g})$. We parameterize g and estimate it from data via maximum marginal likelihood (Eq.˜22) or a suitable surrogate. Finally, we plug the estimate $\hat{\textrm{g}}$ into the conditional posterior (Eq.˜23). Algorithm˜1 presents this recipe.
@@ 3 New Models for Empirical Bayes
We now use BEB to design new methods for empirical Bayes inference. In our first example, we use BEB to recover the classical empirical Bayes g-modeling for the normal means model (§˜3.1). This recovers classical EB methods (Efron2015deconvolution; Soloff2021). We then extend EB to additional probabilistic symmetries: separate/joint exchangeability in arrays (§˜3.2), stationarity in spatial data (§˜3.3), and relative exchangeability conditional on covariates (§˜D.3). In each case, we follow the recipe of Algorithm˜1. We start with a dataset, a symmetry group, and a likelihood. We then design an empirical Bayes model–tailored to the symmetry–and approximate the posterior of the latent structured variables.
@@ 3.1 Classical Example Revisited: Normal Means Model
We assume the true prior $\mathrm{p}^{\star}(\bm{z})$ is exchangeable. By de Finetti’s theorem, we can write Eq.˜24 as the following generative process:
@@
Here $\textrm{g}_{\rm perm}$ is a random prior over $z$. In short, de Finetti’s representation specializes Theorem˜1 to sequences, with $\prod_{i=1}^{\infty}\textrm{g}_{\rm\textrm{perm}}(z_{i})$ the ergodic distribution.
@@
In Bayesian EB, we approximate $\mathrm{p}^{\star}(\textrm{g}_{\rm perm}\mid\bm{x}_{n})$ with its maximum marginal likelihood estimator (MMLE),
@@
If we restrict $\textrm{g}_{\rm perm}$ to a parametric family, then this recovers parametric empirical Bayes estimation, or type-II maximum likelihood. If we consider all possible priors, then it recovers the nonparametric maximum likelihood estimation (NPMLE) approach.
With $\hat{g}_{\rm perm}$ in hand, we approximate the target posterior from Eq.˜26,
@@
The goal is to recover the clean image $\bm{z}^{\star}_{n,n}$ from the observed noisy image $\bm{x}_{n,n}$. The maximum likelihood estimator is the simply the noisy image which has a MSE of $10.291$. The exchangeable sequence approach treats the problem as a simple normal means problem with known variance $\sigma_{0}^{2}$ and applies the NPMLE procedure. The resulting MSE is 3.613.61–a substantial reduction from the error of the observed image. The first three panels in Figure˜1 display the results.
@@ 3.2 Bayesian empirical Bayes for matrix recovery
$u_{i},v_{j},u_{ij}$ $\stackrel{{\scriptstyle iid}}{{\sim}}\textrm{unif}[0,1],\quad\quad\quad i\in[n],j\in[n]$
$z_{ij}$ $=\textrm{g}_{\textrm{sep}}(u_{i},v_{j},u_{ij}),\quad$
$x_{ij}$ $\sim\mathcal{N}(z_{ij},\,\tau_{ij}^{-1}).$
@@
Exact evaluation of Eq.˜30 is intractable because the integral is high–dimensional. We approximate it by maximizing an evidence lower bound (ELBO) under a structured variational family. The estimate $\hat{\textrm{g}}_{\textrm{sep},\theta}$ is the variational MMLE. See Appendix˜D for details.
@@
If the array is jointly exchangeable then it is invariant when rows and columns are permuted together. In this case, Eq.˜28 simplifies to
@@
If the array is relatively exchangeable conditioned on row covariates $\bm{y}_{n}=(y_{i})_{i=1}^{n}$ with $y_{i}\in\mathcal{Y}$ and column covariates $\bm{a}_{p}=(a_{j})_{j=1}^{p}$ with $a_{j}\in\mathcal{A}$, then Eq.˜28 expands to
@@
We revisit the example of denoising an MNIST "5" from Figure˜1. The image is treated as a noisy matrix and we apply the proposed empirical Bayes matrix recovery (EBMR) methods under separate and joint exchangeability. We report $100\times\textrm{MSE}$ between the posterior means $\hat{\bm{z}}_{n,n}$ and the true pixel value $\bm{z}^{\star}_{n,n}$. The fourth and fifth panels in Figure˜1 show the results. EBMR with a jointly exchangeable prior (EBMR–Joint) achieves the lowest error, followed by EBMR with a separately exchangeable prior (EBMR–Sep), then EBMR with a fully exchangeable prior (NPMLE). This ordering aligns with the symmetry hierarchy full $\Rightarrow$ separate $\Rightarrow$ joint: weaker (less restrictive) symmetry assumptions enlarge the class of admissible priors and increase flexibility. On the other hand, weaker symmetry assumptions also increase the sample and computational complexity for estimating g. This creates a trade-off. In our setting, joint exchangeability offers a sweet spot between expressiveness and efficiency.
@@ 3.3 Bayesian empirical Bayes for spatial regression
where $\beta\in\mathbb{R}^{p}$ are regression coefficients and $\bm{z}=(z_{\omega})_{\omega\in\Omega}$ is a latent process that captures spatial dependence. We place a flat prior $\pi(\beta)\propto 1$, so the estimator of $\beta$ coincides with weighted least squares. The resulting model along with Eq.˜32 coincide with the standard model in Bayesian spatial regression (handcock1993bayesian; Stein1999; DeOliveira1997spatial; Berger2001spatial; Banerjee2008; Banerjee2014spatial; Li2024spatial).
@@
$\textrm{g}\sim\mu^{\star},\qquad\bm{z}\mid\textrm{g}\sim\textrm{g},$ (33)
@@
We have defined the invariant variable g as the spectral density of a stationary kernel for a Gaussian process. With this definition, the BEB model is
@@
Figure 1: Denoising a noisy MNIST “5” with Bayesian empirical Bayes methods. Top: original (left), noisy (middle), and NPMLE (right). Bottom: EBMR–sep (left), EBMR–joint (middle), and EB spatial regression (right). Errors are $100\times\textrm{MSE}$ of posterior means; EBMR with a jointly exchangeable prior performs best, followed by EBMR with a separately exchangeable prior and EB spatial regression, then NPMLE.
@@ 4.3 Empirical Bayes Spatial Regression
where the formula for $k_{w,\sigma,\mu}$ is in Eq.˜64. Equivalently, it is a stationary kernel with a three-component Gaussian mixture $0.5\mathcal{N}(0,1^{2})+0.3\mathcal{N}(0.15,0.2^{2})+0.2\mathcal{N}(-0.25,0.35^{2})$ as the spectral density. We draw a single $\bm{z}_{n}^{\star}\sim\mathcal{N}(0,\Sigma^{\star})$ as the ground truth.
@@
Figure˜3 shows the results. EB spatial regression uniformly dominates the regression-adjusted NPMLE, which indicates that stationarity is a more informative symmetry than full exchangeability for this example. At each noise level, the MSE of the EB spatial estimator of $\bm{z}_{n}^{\star}$ decreases sharply with the sample size, while the MSE of the regression-adjusted NPMLE remains essentially flat as $n$ increases. The performance of the two methods is comparable only for small samples (e.g., $n=100$); for larger $n$, the spatial EB estimator achieves substantially lower error. In Figure˜9 of Appendix˜G, we show that the MSE of the posterior-mean estimator for $\beta$ also decreases with $n$ under the EB spatial regression model, which suggests that the posterior of $\beta$ satisfies posterior consistency.
@@ 5.2 Denoising and Interpolating Air Quality Data in NYC
Figure˜5 shows empirical Bayes posterior means with $95\%$ credible bands. The posterior means shows the shrinkage and pooling effects of EB spatial regression: extreme weekly averages (e.g., mid-2023) are pulled toward the mean. Where observations are dense, the posterior uncertainty around the denoised weekly averages is narrow; where data are missing, it is wider. Within Manhattan (Herald Square, Chinatown, Lower East Side), the model borrows heavily across sites—for example, when Chinatown is entirely missing from mid-2021 to early-2022, the model interpolates from Herald Square and the Lower East Side to produce tight bands; similarly, during mid-2023 to early-2024, the model imputes the Lower East Side using the other two Manhattan sites, which recovers a decreasing, cyclical trend. By contrast, Queens North contributes less to interpolating Manhattan gaps (e.g., mid-2022 to early-2023). The reason for this different shrinking behavior is that the EB model finds two spatio–temporal regimes–one for Manhattan and one for northern Queens; see Figure˜10 in §˜G.4.
@@ D.1 EBMR with a Separately Exchangeable Prior
To approximate the intractable marginal likelihood in Eq.˜30, we posit the following structured variational family
@@
Following Eq.˜44, the empirical Bayes estimate minimizes the negative ELBO over the network and grid weights:
@@ Appendix E Algorithms for Empirical Bayes Spatial Regression
where each $\Sigma_{k}=\mathrm{diag}(\sigma_{k,1}^{2},\ldots,\sigma_{k,d}^{2})$ is diagonal. By Bochner’s theorem, Eq.˜63 induces the kernel
@@ G.2 Covariate-Assisted Empirical Bayes with Model Misspecification
G.3 Empirical Bayes Spatial Regression Results
Figure˜9 shows that the MSE of the posterior mean estimator of $\beta$ relative to $\beta^{\star}$ decreases as the sample size $n$ grows, for different choices of the number of spectral mixture components $K$.
G.4 Additional Results for NYC Air Quality